Calculate the pH of each solution at 25 ∘C .

1.0×10−3 M HClpH=

0.01 M HNO3pH=

1.0×10−3 M NaOHpH=

0.1 M KOHpH=
Answer #1

1)

[H+] = [HCl] = 1.0*10^-3 M

use:

pH = -log [H+]

= -log (1*10^-3)

= 3

Answer: 3.00

2)

[H] = [HNO3] = 0.01 M

use:

pH = -log [H+]

= -log (1*10^-2)

= 2

Answer: 2.00

3)

[OH-] = [NaOH] = 1.0*10^-3 M

use:

pOH = -log [OH-]

= -log (1*10^-3)

= 3

use:

PH = 14 - pOH

= 14 - 3

= 11

Answer: 11.00

4)

[OH-] = [KOH] = 0.1 M

use:

pOH = -log [OH-]

= -log (0.1)

= 1

use:

PH = 14 - pOH

= 14 - 1

= 13

Answer: 13.00